The length of a retangle is five less than twice the width. The perimeter is 50 cm. Using an equation,find the dememsions of the rectangle.
Answer by anantha(86)

sol:
let the width be x
twice the width=2x
length=5 less than the twice the width
length=2x-5
perimeter of the rectangle=2*(length+width)
=2*(2x-5+x)
=2*(3x-5)
=2*3x-2*5
=6x-10
according to the problem perimeter=50 cm
6x-10=50
adding 10 on both sides
6x-10+10=50+10
6x=60
dividing by 6 on both sides
6x/6=60/6
x=10
width=10 cm
length=2x-5
=2*10-5
=20-5
=15 cm
length=15cm,width=10cm